Paula plans to replace the vinyl floor covering her 10-ft by 12-ft kitchen. She wants to have a border of even width of a special material. She can afford only 21 ft^2
of this material. How wide a border can she have?

Answer by Boreal(15235)

Draw this.
let the border=x
the covering will be 10-2x by 12-2x, because the border affects both sides.
that area is 21 ft^2,
so 120-44x+4x^2=21
4x^2-44x+99=0
x=(1/8)(44+/- sqrt (1936-1584)); sqrt term is 18.76
x=25.24/8 the other x would be nearly 8 and wouldn't work.
x=3.155 feet border, 6.310 feet taken from both sides.
the floor would be 10-6.310 or3.69 by 12-6.310 or 5.69
3.69*5.69=20.9961
The answer is 3.155 feet.